OpenStudy (anonymous):

A drink costs $1 plus half its price, how much does it cost?

OpenStudy (anonymous):

1.50

OpenStudy (anonymous):

lol it aint that easy , its a good ques

OpenStudy (anonymous):

think about it deeply, dont just give me an ans

OpenStudy (saifoo.khan):

\[\Huge $1.50only\]

OpenStudy (anonymous):

1+ 1/2 C=C

OpenStudy (anonymous):

C being coost

OpenStudy (anonymous):

you all are wrong, lets see who answers this one

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

i thought it was 1.50 and i asked this question to my best friend google and he gave me a link to a similar problem let the cost of a drink be x drink costs $1 plus half its price p=1+(p/2) p=(2+p)/2 2p=2+p 2p-p=2 so,p=2 this is a pretty cool problem :D

OpenStudy (anonymous):

oh true, its a really cool one ;)

OpenStudy (anonymous):

that's the answer posted :p But yeah, it's still a cool problem :) (did you mean p=price? U said x)

OpenStudy (anonymous):

oh sorry yes

OpenStudy (anonymous):

too expensive for me :)